Find the equation of the line passing through the point (3,2,-8) and is perpendicular to the plane     -3x+y+2z-2=0. 
Book answer:(x-3)/3=(y+2)/1=(z+8)/2
 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
Answer should be
 
(x- 3)/ – 3 = (y- 2) /1 = (z+8)/2
 
Answer , which you have written above is incorrect.
Please check and let me know